During rush hour, Adriana can drive 20 miles using the side roads in the same time that it takes to travel 15 miles on the freeway. If Adriana's rate on the side roads is 10 mi/h faster than her rate on the freeway, find her rate on the side roads.
Thank You

Answer by stanbon(48553)

During rush hour, Adriana can drive 20 miles using the side roads in the same time that it takes to travel 15 miles on the freeway. If Adriana's rate on the side roads is 10 mi/h faster than her rate on the freeway, find her rate on the side roads.
----------------------------------------
Side road DATA:
distance= 20 mi ; rate= x mph ; time= d/r= 20/x hrs.
------------------------------
Freeway DATA:
distance= 15 mi ; rate=x-10 mph; time= d/r = 15/(x-10)
--------------------------------
EQUATION:
time = time
20/x = 15/(x-10)
20x-200=15x
5x=200
x=40 mph (Her rate on the side road)
Cheers,
Stan H.